OpenStudy (anonymous):

Solve the following problem for the roots by completing the square. 3x2 -12x = 9

OpenStudy (anonymous):

Still need help with this or did you get it already?

OpenStudy (anonymous):

@JoyLin I still need help please.

OpenStudy (anonymous):

divide by 3 first

OpenStudy (anonymous):

6x-12x=9 6x-12x=-6x -6x=9 /-6 /-6 gives you the answer